OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle is 2 cm. Find the measure of one of its legs.

OpenStudy (saifoo.khan):

measure?

OpenStudy (anonymous):

in cm

OpenStudy (anonymous):

\[\sqrt{2}\]

OpenStudy (anonymous):

a^2 + b^2 = 2^2, a = b a = b = sqrt(2)

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] in this case, the hypotenuse is c, and the legs are the same as each other, so:\[2a^2=4\]\[a^2=2\]\[a=\sqrt{2}\]